# Adaptive Querying for Reward Learning from Human Feedback

### v1

@@ Abstract
Learning from human feedback is a popular approach to train robots to adapt to user preferences and improve safety. Existing approaches typically consider a single querying (interaction) format when seeking human feedback and do not leverage multiple modes of user interaction with a robot. We examine how to learn a penalty function associated with unsafe behaviors, such as side effects, using multiple forms of human feedback, by optimizing the query state and feedback format. Our framework for adaptive feedback selection enables querying for feedback in critical states in the most informative format, while accounting for the cost and probability of receiving feedback in a certain format. We employ an iterative, two-phase approach which first selects critical states for querying, and then uses information gain to select a feedback format for querying across the sampled critical states. Our evaluation in simulation demonstrates the sample efficiency of our approach.
@@ I INTRODUCTION
A key factor affecting an autonomous agent’s behavior is its reward function. Due to the complexity of real-world environments and the practical challenges in reward design, agents often operate with incomplete reward functions corresponding to underspecified objectives, which can lead to unintended and undesirable behaviors such as negative side effects (NSEs) [1, 2, 3]. For example, an indoor robot that optimizes distance to goal may break a vase as a side effect if its reward function does not model the undesirability of breaking a vase while navigating to the goal [4] (Figure 1).
@@
Several prior works have examined learning from various forms of human feedback to improve robot performance, including avoiding side effects [5, 6, 7, 8, 9, 10, 11]. In many real-world settings, the human can provide feedback in many forms, ranging from binary signals indicating action approval to correcting robot actions, each varying in the granularity of information revealed to the robot and the human effort required to provide it. To efficiently balance the trade-off between seeking feedback in a format that accelerates robot learning and reducing human effort involved, it is beneficial to seek detailed feedback sparingly in certain states and complement it with feedback types that require less human effort in other states. Such an approach could also reduce the sampling biases associated with learning from any one format, thereby improving learning performance [12]. In fact, a recent study indicates that users are generally willing to engage with the robot in more than one feedback format [13]. Existing approaches utilize a single feedback format throughout the learning process and do not support gathering feedback in different formats in different regions of the state space [14, 15].
How can a robot identify when to query and in what format, while accounting for the cost and availability of different forms of feedback? We present a framework for adaptive feedback selection (AFS) that enables a robot to seek feedback in multiple formats in its learning phase, such that its information gain is maximized. In the interest of clarity, AFS is introduced in the context of NSEs but the framework is general and can be applied broadly.
Figure 1: An illustration of adaptive feedback selection. Agent learns to navigate to the goal without breaking vases, by querying the human in different format across the state space. Red circles show critical states along with queries, and orange speech bubbles show user feedback.
The information gain of a feedback format is measured as the Kullback–Leibler (KL) divergence between the true NSE distribution, revealed to the robot via human feedback collected so far, and the robot’s current knowledge of NSEs based on the feedback it has received. In each querying cycle, the robot selects a feedback format that maximizes its information gain, given its current knowledge of NSEs.
When collecting feedback in every state is infeasible, the robot must prioritize querying in critical states—states where human feedback is crucial for learning an association of state features and NSEs, i.e., a predictive model of NSE severity. Querying in critical states maximizes information gain about NSEs, compared to other states. Prior works, however, query for feedback in states randomly sampled or along the shortest path to the goal, which may not result in a faithful NSE model [2, 11].
@@
We use an iterative approach to gather NSE information under a limited query budget (Figure 2). The key steps are: (1) states are partitioned into clusters, with a cluster weight proportional to the number of NSEs discovered in it; (2) a critical states set is formed by sampling from each cluster based on its weight; (3) a feedback format that maximizes the information gain in critical states is identified, while accounting for the cost and uncertainty in receiving a feedback, using the human feedback preference model; and (4) cluster weights and information gain are updated, and a new set of critical states are sampled to learn about NSEs, until the querying budget expires. The learned NSE information is mapped to a penalty function and augmented to the robot’s model to compute an NSE-minimizing policy to complete its task. Empirical evaluation on four domains in simulation demonstrate the effectiveness of our approach in learning to mitigate NSEs from explicit and implicit feedback types.
II BACKGROUND
@@
Markov Decision Processes (MDPs) are a popular framework to model sequential decision making problems. An MDP is defined by the tuple $M\!=\!\langle S,A,T,R,\gamma\rangle$, where $S$ is the set of states, $A$ is the set of actions, $T(s,a,s^{\prime})$ is the probability of reaching state $s^{\prime}\!\in\!S$ after taking an action $a\!\in\!A$ from a state $s\!\in\!S$ and $R(s,a)$ is the reward for taking action $a$ in state $s$. An optimal deterministic policy $\pi^{*}:\!S\!\rightarrow\!A$ is one that maximizes the expected reward. When the objective or reward function is incomplete, even an optimal policy can produce unsafe behaviors such as side effects.
Negative Side Effects (NSEs) are immediate, undesired, unmodeled effects of an agent’s actions on the environment [16, 17, 3]. We focus on NSEs arising due to incomplete reward function [2], which we mitigate by learning a penalty function using human feedback.
Learning from Human Feedback is a widely used technique to train agents when reward functions are unavailable or incomplete [18, 9, 19], including to improve safety [20, 21, 22, 23, 11, 2]. Feedback can take various forms such as demonstrations [24, 25], corrections [26, 27, 28], critiques [5, 2], ranking trajectories [29], or may be implicit in the form of facial expressions and gestures [6, 30]. Existing approaches focus on learning from a single feedback type, limiting learning efficiency. Recent studies consider combinations such as demonstrations and preferences [31, 32], but assume a fixed order and do not scale to multiple formats. Another recent work examines feedback format selection by estimating the human’s ability to provide feedback in a certain format [33]. Unlike these approaches, we dynamically select the most informative feedback without any pre-processing.
The information gain associated with a feedback quantifies the effect of a feedback in improving the agent’s understanding of the underlying reward function, often measured using Kullback-Leibler (KL) Divergence [33, 34], $D_{KL}(P\|Q)=\sum_{x}P(x)\log\frac{P(x)}{Q{x}}$ where $P$ is the prior distribution and $Q$ is the posterior distribution after observing evidence.
@@ III PROBLEM FORMULATION
Setting: Consider a robot operating in an environment modeled as a Markov Decision Process (MDP), using its acquired model $M=\langle S,A,T,R_{T}\rangle$. The robot optimizes the completion of its assigned task, which is its primary objective described by reward $R_{T}$. A primary policy, $\pi^{M}$, is an optimal policy for the robot’s primary objective.
@@
This work assumes the robot has access to the user’s feedback preference model $D$—either handcrafted by an expert or learned from user interactions prior to robot querying. Abstracting user feedback preferences into probabilities and costs enables generalizing the preferences across similar tasks. We take the pragmatic stance that $\psi$ is independent of time and state, denoting the user’s preference about a format, such as not preferring formats that require constant supervision of robot performance. While this can be relaxed and the approach can be extended to account for state-dependent preferences, getting an accurate state-dependent $\psi$ could be challenging in practice.
@@
Below, we describe how a robot can learn a penalty function associated with NSEs ($\hat{R}_{N}$), given data from different types of feedback. In Section IV, we describe when and how to seek feedback, given feedback preference model $D$.
III-A Learning $\hat{R}_{N}$ from multiple forms of feedback
Since the agent has no prior knowledge about NSEs, it assumes none of its actions produce NSEs. We examine learning an NSE penalty function $\hat{R}_{N}$ using the following popular feedback formats and their annotated (richer) versions. In our settings, an action in a state may cause either mild, severe, or no NSEs. In practice, any number of NSE categories can be considered, provided the feedback formats align with them.
@@ III-A Learning $\hat{R}_{N}$ from multiple forms of feedback
Gaze:  In this implicit feedback format, the robot requests to collect gaze data of the user and compares its action outcomes with the gaze positions of the user [10]. Actions with outcomes aligning with the average gaze direction are labeled as acceptable, and unacceptable otherwise.
Each of the above format provides different levels of detail, thereby resulting in different learned reward models. Figure 3 shows the interaction format and learned reward values, using different feedback types in isolation, on the vase domain. Breaking a vase on a carpet is a mild NSE. Breaking a vase on a hard surface is a severe NSE.
NSE Model Learning: We use $l_{m}$, $l_{h}$, and $l_{a}$ to denote labels corresponding to mild, severe and no NSEs respectively. An acceptable action in a state is mapped to label $l_{a}$, $(s,a)\rightarrow l_{a}$, while unacceptable action is mapped to label $l_{h}$. If the NSE severity of unacceptable actions are known, then actions with mild NSE are mapped to $l_{m}$ and those with severe NSEs mapped to $l_{h}$. Mapping feedback to these labels provides a consistent representation of NSE severity for learning under various feedback types. The NSE severity labels, derived from the gathered feedback, are generalized to unseen states by training a random forest classifier (RF) model to predict NSE severity of an action in a state. Any classifier can be used in practice. Hyperparameters for training are determined by a randomized search in the RF parameter space, using three-fold cross validation and selecting parameters with the least mean squared error for training and subsequently, determining the NSE severity. The label for each state-action pair is then mapped to its corresponding penalty value, yielding $\hat{R}_{N}(s,a)$. In our experiments, the penalties for $l_{a}$, $l_{m}$, and $l_{h}$ are 00, $+5$, and $+10$ respectively.
IV ADAPTIVE FEEDBACK SELECTION
Given an agent’s decision making model $M$ and the human’s feedback preference model $D$, adaptive feedback selection (AFS) enables the agent to query for feedback in critical states in a format that maximizes its information gain.
Let $p^{*}$ be the true underlying NSE distribution, unknown to the agent but known to the human. $p^{*}$ deterministically maps state-action pairs to an NSE severity level (i.e., no NSE, mild NSE or severe NSE). Human feedback, when available, is sampled from this distribution. Let $p\sim p^{*}$ denote the distribution of state-action pairs causing NSE of varying severity, aggregated from all feedback received thus far. In other words, $p$ represents the accumulated NSE information known to the agent, based on human feedback. Let $q$ denote the agent’s learned NSE distribution, based on all feedback received up to that point i.e., from $p$.
Below we describe an approach to select critical states, followed by an approach for feedback format selection, based on the KL divergence between $p$ and $q$.
IV-A Critical States Selection
Intuitively, when the budget for querying a human is limited, it is useful to query in states with a high learning gap—the divergence between the agent’s knowledge of NSE distribution and the underlying NSE distribution, given feedback data collected so far. States with a high learning gap are called critical states ($\Omega$) and querying in these states can reduce the learning gap. The learning gap at iteration $t$ is measured as the KL divergence between the information gathered so far ($p^{t}$) and the agent’s learned NSE distribution ($q^{t-1}$): $D_{KL}(p^{t}\|q^{t-1})$.
We compare $p^{t}$ with $q^{t-1}$ since we want to identify states where the agent’s knowledge of NSE was incorrect, thereby guiding the selection of next batch of critical states. While $D_{KL}(p^{t}\|q^{t})$ may seem to be a better choice to guide critical states selection, this measure only shows how well the agent learned using the feedback at $t$ but does not reveal states where the agent was incorrect about NSEs. Algorithm 1 outlines our approach for selecting critical states at each learning iteration, with the following three key steps.
2. Estimating information gain: We define the information gain of sampling from a cluster $k\!\in\!K$, based on the learning gap discussed earlier,
$IG(k)^{t}\!$ $=\!\frac{1}{|\Omega_{k}^{t-1}|}\sum_{s\in\Omega_{k}^{t-1}}D_{KL}(p^{t}\|q^{t-1})$ (1)
$=\!\frac{1}{|\Omega_{k}^{t-1}|}\sum_{s\in\Omega_{k}^{t-1}}\sum_{a\in A}p^{t}(a|s)\cdot\log\left(\frac{p^{t}(a|s)}{q^{t-1}(a|s)}\right)$ (2)
where $\Omega_{k}^{t-1}$ denotes the set of states sampled for querying from cluster $k$ at iteration $t-1$. $p^{t}(a|s)$ and $q^{t-1}(a|s)$ denote the NSE severity labels of action $a$ in $s$, as provided in the feedback data and in the agent’s learned model, respectively.
3. Sampling critical states: At each learning iteration $t$, the agent assigns a weight $w_{k}$ to each cluster $k\!\in\!K$, proportional to the new information on NSEs revealed by the most informative feedback format identified at $t-1$, using Eqn. 2. Clusters are given equal weights when there is no prior feedback (Line 4). We sample critical states in batches but they can also be sampled sequentially. When sampling in batches of $N$ states, the number of states $n_{k}$ to be sampled from each cluster is determined by its assigned weight. At least one state is sampled from each cluster to ensure sufficient information for calculating the information gain for every cluster (Line 5). The agent randomly samples $n_{k}$ states from corresponding cluster and adds them to a set of critical states $\Omega$ (Lines 6, 7). If the total number of critical states sampled is less than $N$ due to rounding, then the remaining $N_{r}$ states are sampled from the cluster with the highest weight and added to $\Omega$ (Lines 9-12).
Algorithm 1 Critical States Selection
0:  $N$: #critical states; $\mathcal{K}$:#clusters
1:  $\Omega\leftarrow\emptyset$
@@ IV-A Critical States Selection
4:     $W_{k}\leftarrow\begin{cases}\frac{1}{\mathcal{K}},\text{ if no feedback received in any iteration}\\ \frac{IG(k)}{\sum_{k\in K}IG(k)},\text{ if feedback received}\end{cases}$

### v2

@@ Abstract
Learning from human feedback is a popular approach to train robots to adapt to user preferences and improve safety. Existing approaches typically consider a single querying (interaction) format when seeking human feedback and do not leverage multiple modes of user interaction with a robot. We examine how to learn a penalty function associated with unsafe behaviors using multiple forms of human feedback, by optimizing both the query state and feedback format. Our proposed adaptive feedback selection is an iterative, two-phase approach which first selects critical states for querying, and then uses information gain to select a feedback format for querying across the sampled critical states. The feedback format selection also accounts for the cost and probability of receiving feedback in a certain format. Our experiments in simulation demonstrate the sample efficiency of our approach in learning to avoid undesirable behaviors. The results of our user study with a physical robot highlight the practicality and effectiveness of adaptive feedback selection in seeking informative, user-aligned feedback that accelerate learning. Experiment videos, code and appendices are found on our website: https://tinyurl.com/AFS-learning
@@ I INTRODUCTION
A key factor affecting an autonomous agent’s behavior is its reward function. Due to the complexity of real-world environments and the practical challenges in reward design, agents often operate with incomplete reward functions corresponding to underspecified objectives, which can lead to unintended and undesirable behaviors such as negative side effects (NSEs) [2, 35, 43]. For example, a robot optimizing the distance to transport an object to the goal, may damage items along the way if its reward function does not model the undesirability of colliding into other objects in the way (Figure 1).
Figure 1: An illustration of adaptive feedback selection. The robot arm learns to move the blue object to the white bin, without colliding with other objects in the way, by querying the human in different format across the state space.
@@
Human feedback offers a natural way to provide the missing knowledge, and several prior works have examined learning from various forms of human feedback to improve robot performance, including avoiding side effects [13, 14, 28, 31, 39, 52]. In many real-world settings, the human can provide feedback in many forms, ranging from binary signals indicating action approval to correcting robot actions, each varying in the granularity of information revealed to the robot and the human effort required to provide it. For instance, a person supervising a household robot may occasionally be willing to provide detailed corrections when the robot encounters a fragile vase but may only want to give quick binary approvals during a routine motion. Ignoring this variability either limits what the robot can learn or burdens the user. To efficiently balance the trade-off between seeking feedback in a format that accelerates robot learning and reducing human effort involved, it is beneficial to seek detailed feedback sparingly in certain states and complement it with feedback types that require less human effort in other states. Such an approach could also reduce the sampling biases associated with learning from any one format, thereby improving learning performance [36]. In fact, a recent study indicates that users are generally willing to engage with the robot in more than one feedback format [37]. However, existing approaches rarely exploit this flexibility, and do not support gathering feedback in different formats in different regions of the state space [12, 41].
These practical considerations motivate the core question of this paper: “How can a robot identify when to query and in what format, while accounting for the cost and availability of different forms of feedback?” We present a framework for adaptive feedback selection (AFS) that enables a robot to seek feedback in multiple formats in its learning phase, such that its information gain is maximized. Rather than treating all states and feedback formats uniformly, AFS prioritizes human feedback in states where feedback is most valuable and chooses feedback types based on their expected cost and information gain. This design reduces user effort, accommodates different levels of feedback granularity, and focuses on states where learning contributes to safety. In the interest of clarity, the rest of this paper grounds the discussion of AFS as an approach for robots to learn to avoid negative side effects (NSEs) of their actions. The NSEs refer to unintended and undesirable outcomes that arise as the agent performs its assigned task. In object delivery example in Figure 1, the robot may inadvertently collide with other objects on the table, producing NSEs. Focusing on NSEs provides a well-defined and measurable setting–quantified by the number of NSE occurrences–to evaluate how AFS improves an agent’s learning efficiency and safety. However, note that AFS is a general technique that can be applied broadly to learn about various forms of undesirable behavior.
@@
Minimizing NSEs using AFS involves four iterative steps (Figure 4): (1) states are partitioned into clusters, with a cluster weight proportional to the number of NSEs discovered in it; (2) a set of critical states—states where human feedback is crucial for learning an association of state features and NSEs, i.e., a predictive model of NSE severity, is formed by sampling from each cluster based on its weight; (3) a feedback format that maximizes the information gain in critical states is identified, while accounting for the cost and uncertainty in receiving a feedback, using the human feedback preference model; and (4) update cluster weights and information gain, and sample a new set of critical states to learn about NSEs, until the querying budget expires. The learned NSE information is mapped to a penalty function and augmented to the robot’s model to compute an NSE-minimizing policy to complete its task.
We evaluate AFS in both simulation and using a user study where participants interact with a robot arm. First, we evaluate the approach in three simulated proof-of-concept settings with simulated human feedback. Second, we conduct a pilot study where 12 human participants interact with and provide feedback to the agent in a simulated gridworld domain. Finally, we evaluate using a Kinova Gen3 7DoF arm and 30 human participants. Besides the performance and sample efficiency, our experiments also provide insights into how the querying process can influence user trust. Together, these complementary studies demonstrate both the practicality and effectiveness of AFS.
II BACKGROUND and RELATED WORK
II-A Markov Decision Processes (MDPs)
@@
The MDPs are a popular framework to model sequential decision making problems. An MDP is defined by the tuple $M\!=\!\langle S,A,T,R,\gamma\rangle$, where $S$ is the set of states, $A$ is the set of actions, $T(s,a,s^{\prime})$ is the probability of reaching state $s^{\prime}\!\in\!S$ after taking an action $a\!\in\!A$ from a state $s\!\in\!S$ and $R(s,a)$ is the reward for taking action $a$ in state $s$. An optimal deterministic policy $\pi^{*}:\!S\!\rightarrow\!A$ is one that maximizes the expected reward. When the objective or reward function is incomplete, even an optimal policy can produce unsafe behaviors such as side effects. Negative Side Effects (NSEs) are immediate, undesired, unmodeled effects of an agent’s actions on the environment [26, 38, 43]. We focus on NSEs arising due to incomplete reward function [35], which we mitigate by learning a penalty function using human feedback.
II-B Learning from Human Feedback
Learning from human feedback is a popular approach to train agents when reward functions are unavailable or incomplete [1, 31, 34, 30], including to improve safety [7, 8, 17, 33, 52, 35, 19]. Feedback can take various forms such as demonstrations [32, 35, 40, 50], corrections [11, 3], critiques [13, 45], ranking trajectories [6, 47, 15], natural language instructions [29, 48, 19] or may be implicit in the form of facial expressions and gestures [14, 44, 9].
While the existing approaches for learning from feedback have shown success, they typically assume that a single feedback type is used to teach the agent. This assumption limits learning efficiency and adaptability. Some efforts combine demonstrations with preferences [5, 21], showing that utilizing more than one format accelerates learning. Extending this idea, recent works integrate richer modalities such as language and vision with demonstrations.  [49] learn reward function from comparative language feedback, while  [42] show that a single demonstration or natural language description can help define a proxy reward when used along with a vision-language models (VLM) that is pretrained on a large amount of out-of-domain video demonstrations and language pairs.  [24] use multimodal embeddings of visual observations and natural language descriptions to compute alignment-based rewards. A recent study even emphasizes that combining multiple feedback modalities can further enhance learning outcomes [4]. Together, these works highlight that combining complementary feedback formats help advance reward learning beyond using a fixed feedback format. In contrast, our approach uses multiple forms of human feedback for learning.
Other approaches that learn from human feedback focus on modeling variations in human behavior. [20] model the heterogeneous behaviors of human, capturing differences in feedback frequency, delay, strictness, and bias to improve the robustness during the learning process, as optimal behaviors vary across users. Along the same line, the reward learning approach proposed by [16], selects a single feedback format based on the user ability to provide feedback in that format, resulting in an interaction that is tailored to a user’s skill level. Collectively, these works reveal a shift towards adaptive and user-aware querying mechanisms that improve reward inference and learning efficiency, motivating our approach to dynamically select both when to query and in what feedback format.
@@ III PROBLEM FORMULATION
Setting: Consider a robot operating in a discrete environment modeled as a Markov Decision Process (MDP), using its acquired model $M=\langle S,A,T,R_{T}\rangle$. The robot optimizes the completion of its assigned task, which is its primary objective described by reward $R_{T}$. A primary policy, $\pi^{M}$, is an optimal policy for the robot’s primary objective.
@@
Running Example: We illustrate the problem using a simple object delivery task using a Kinova Gen3 7DoF arm shown in Figure 1. The robot optimizes delivering the blue block to the white bin, by taking the shortest path. However, passing through states with a cardboard box or a glass bowl constitutes an NSE. Since the robot has no prior knowledge about NSEs of its actions, it may inadvertently navigate through these states causing NSEs.
@@
This work assumes the robot has access to the user’s feedback preference model DD—either handcrafted by an expert or learned from user interactions prior to robot querying, as in our user study experiments. Abstracting user feedback preferences into probabilities and costs enables generalizing the preferences across similar tasks. We take the pragmatic stance that $\psi$ is independent of time and state, denoting the user’s preference about a format, such as not preferring formats that require constant supervision of robot performance. While this can be relaxed and the approach can be extended to account for state-dependent preferences, obtaining an accurate state-dependent $\psi$ could be challenging in practice.
@@
Below, we describe the various feedback formats considered in this paper, and how the data from these formats are mapped to NSE severity labels.
III-A Feedback Formats Studied
The agent learns an association between state-action pairs and NSE severity, based on the human feedback provided in response to agent queries. The NSE categories we consider in this work are $\{\text{No NSE},\text{Mild NSE},\text{Severe NSE}\}$. We focus on the following commonly used feedback types, each differing in the level of information conveyed to the agent and the human effort required to provide them.
@@ III-A Feedback Formats Studied
Rank:  The robot randomly selects $N$ ranking queries of the form $\langle\textit{state},\textit{action 1},\textit{action 2}\rangle$, by sampling two actions for each critical state. The human selects the safer action among the two options. If both are safe or unsafe, one of them is selected at random. The selected action is marked as acceptable and the other is treated as unacceptable.
Mapping feedback data to NSE severity labels: We use $l_{a}$, $l_{m}$, and $l_{h}$ to denote labels corresponding to no, mild and severe NSEs, respectively. An acceptable action in a state is mapped to $l_{a}$, i.e., $(s,a)\rightarrow l_{a}$, while an unacceptable action is mapped to $l_{h}$. When the severity of NSEs for unacceptable actions is known, actions producing mild NSEs are mapped to $l_{m}$ and those producing severe NSEs to $l_{h}$. Mapping feedback to this common label set provides a consistent representation of NSE severity across diverse feedback types. The granularity of information and the sampling biases of the different feedback types affect the learned reward. Figure 2 illustrates this with the learned NSE penalty for the running example of moving an object to the bin (Fig. 1), motivating the need for an adaptive approach that can learn from more than one feedback format. In the running example, the robot arm colliding with cardboard boxes is a mild NSE, and colliding with a glass bowl is a severe NSE.
IV ADAPTIVE FEEDBACK SELECTION
Given an agent’s decision making model $M$ and the human’s feedback preference model $D$, AFS enables the agent to query for feedback in critical states in a format that maximizes its information gain. We first formalize the NSE model learning process and then describe in detail how AFS selects critical states and the query format.
Formalizing NSE Model Learning: Let $p^{*}:S\times A\rightarrow\{l_{a},l_{m},l_{h}\}$ denote the true NSE severity label for each state-action pair, which is unknown to the agent but known to the human. The label $l_{a}$ corresponds to no NSE, $l_{m}$ denotes mild NSE, $l_{h}$ denote the label for severe NSE. Let $p$ be a sampled approximation of $p^{*}$ ($p\sim p^{*}$), denoting the dataset of NSE labels collected via human feedback in response to the $(s,a)$ pairs queried. That is, $p^{t}$ denotes the data collected from human feedback until iteration $t$, where $p^{t}(s,a)$ represents the categorical NSE severity label assigned to the state-action pair $(s,a)$. Let $q:S\times A\rightarrow\{l_{a},l_{m},l_{h}\}$ denote the labels predicted by the learned NSE model—learned using a supervised classifier with $p$ as the training data. In this paper, we use a Random Forest (RF) classifier, though any classifier can be used in practice. Hyperparameters are optimized through randomized search with three-fold cross validation, and the configuration yielding the lowest mean-squared error is selected for training.
Figure 3 shows an example of $p$ and $q$ for the object delivery task. We encode NSE categories as $\{0,1,2\}$ corresponding to $\{$ no NSE, mild NSE, severe NSE}\} respectively. Each state has four possible actions $A\!=\!\{a_{1},a_{2},a_{3},a_{4}\}$, and the vector $p(s)\!=\!\{\cdot,\cdot,\cdot,\cdot\}$ (and similarly $q(s)$) encodes the categorical NSE labels for $(s,a_{1}),(s,a_{2}),(s,a_{3}),(s,a_{4})$ in that order. Since the human’s categorization of NSE is initially unknown, $p(s)$ is sampled from a uniform prior over the labels, and $q(s)$ is initialized to $[0,0,0,0]$ (all actions are assumed to be safe) across all states.
At t1t-1, $p^{t-1}$ reflects a single labeled state from the feedback received, while $q^{t-1}$ reflects NSE label for the state after learning from $p^{t-1}$. For example, in iteration t1t\!-\!1, an action $a_{3}$ in state $s$ is randomly selected for querying using the Annotated Approval feedback format. The human labels it as mild NSE, so $p^{t-1}(s,a_{3})=1$, and consequently $p^{t-1}(s)=[0,0,1,0]$. After training on $p^{t-1}$, the classifier may sometimes incorrectly predict $q^{t-1}(s)\!=\![0,0,0,0]$, especially in early iterations when there is less data. At the next iteration $t$, the agent queries in a similar state using the Approval format, where the action $a_{1}$ is randomly selected. Because the NSE severity level (i.e., mild/severe) cannot be indicated through the Approval format, $p^{t}$ is updated as $p^{t}(s)=[2,0,0,0]$, and training now yields a prediction $q^{t}(s)=[2,0,1,0]$ (i.e. the NSE model predicts severe NSE outcome on $a_{1}$ and a mild NSE outcome on $a_{3}$). This illustrates that $q$ may initially disagree with $p$, but as feedback accumulates on related states, the generalization of $q$ across actions begins to align with $p$.
Each predicted label is then mapped to a penalty value to form the learned penalty function, $\hat{R}_{N}$, with penalties for $l_{a},l_{m}$ and $l_{h}$ set to 0,50,-5 and 10-10 respectively, in our experiments. This penalty function is integrated into the agent’s reward model to compute an updated policy that minimizes NSEs while completing the primary task.
Figure 3: Illustration of $p$ (accumulated feedback) and $q$ (generalized NSE labels) for the object delivery task. $f^{*}_{1:t-1}$ indicates the feedback formats selected until iteration t1t-1. Refer to caption indicates no NSE; Refer to caption indicates mild NSE; Refer to caption indicates severe NSE. Queried states in each iteration is highlighted in blue.
In this learning setup, minimizing NSEs using AFS involves four iterative steps (Figure 4). In each learning iteration, AFS identifies (1) which states are most critical for querying (Sec. IV-A), and (2) which feedback format maximizes the expected information gain at the critical states, while accounting for user feedback preferences and effort involved (Sec. IV-B). The information gain associated with a feedback quantifies the effect of a feedback in improving the agent’s understanding of the underlying reward function, and is measured using Kullback-Leibler (KL) Divergence [16, 46]. At the end of each iteration, the cluster weights and information gain are updated, and a new set of critical states are sampled to learn about NSEs, until the querying budget expires or the KL-divergence is below a problem-specific, pre-defined threshold.
IV-A Critical States Selection
When the budget for querying a human is limited, it is useful to query in states with a high learning gap measured as the KL-divergence between the agent’s knowledge of NSE severity and the true NSE severity given the feedback data collected so far. States with a high learning gap are called critical states ($\Omega$) and querying in these states can reduce the learning gap.
Algorithm 1 Critical States Selection
0:NN: #critical states; $\mathcal{K}$:#clusters;
@@ IV-A Critical States Selection
[…]